Quantitative Problem: Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis. Both projects' after-tax cash flows are shown on the time line below. Depreciation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firm's average project. Bellinger's WACC is 9%.

0 1 2 3 4
Project A -1,350 700 425 290 340
Project B -1,350 300 360 440 790

What is Project A's MIRR? Do not round intermediate calculations. Round your answer to two decimal places.

________%

What is Project B's MIRR? Do not round intermediate calculations. Round your answer to two decimal places.

________%
Answer #1

A:

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value of inflows=700*(1.09)^3+425*(1.09)^2+290*(1.09)+340

=2067.5628

MIRR=[Future value of inflows/Present value of outflows]^(1/time period)=1

=[2067.5628/1350]^(1/4)-1

=11.25%(Approx).

B:

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value of inflows=300*(1.09)^3+360*(1.09)^2+440*(1.09)+790

=2085.8247

MIRR=[Future value of inflows/Present value of outflows]^(1/time period)=1

=[2085.8247/1350]^(1/4)-1

=11.49%(Approx).